### v1

# Self-organisation - the underlying principle and a general model

Abstract
Recent observations of coordinated self-organisation (SO) of stress and structure in granular systems provide insight into the fundamental principle underlying this phenomenon. It is first argued here that SO emerges when a minute subset of configurations are significantly more stable than the rest and therefore survive the noise in the system much longer to be observed. This principle goes deeper than recently proposed energy considerations. Guided by this principle, a statistical mechanics model is formulated then for SO in these systems and its extension to three dimensions is outlined. The principle holds beyond granular systems and the model is extended next to describe emergence of SO in more general systems. The application of the model is illustrated for the specific example of laning. Parallels of the modelling approach to traditional statistical mechanics provide useful insight that should assist in modelling SO in other out-of-equilibrium systems.
I. Introduction
The term self-organisation (SO) refers usually to spontaneous emergence of large-scale ordered patterns or symmetries in driven disordered out-of-equilibrium systems. This phenomenon is ubiquitous, from formation of structural patterns in inanimate matter, such as snowflakes, dune ripples, and some cloud formations, to temporal manifestations and evolution of living organisms. SO, often associated with complexity and pattern formation, emerges as a result of local interactions between the individual system constituents and is not externally engineered. While this description also applies to traditional phase transitions in thermal systems, in which the competition is between energy and entropy, it is much more general, manifesting in physical, chemical, biological, active, and social systems SOR1 ; SOR2 ; SOR3 ; SOR4 ; SOR5 .
Most studies of SO have been usually focused on specific contexts and in many case are descriptive. Attempts at finding a general principle underpinning SO are few and they are mainly based on energy considerations. For example, it has been proposed that SO emerges in steady states by selection of microstates that minimise energy-dissipation in driven systems Cyetal21 ; Keetal23 . These models propose that the system random-walks in the multi-dimensional microstates space and those special microstates are selected as the system self-organises. However, this does not explain what is exactly the physical mechanism that makes the random walk park on the microstates of low energy dissipation. Moreover, this idea excludes emergence of SO in out-of-equilibrium a-thermal systems, where energy plays a negligible role. Observations of SO in these systems suggest that a more general principle is in play, which applies to all systems, however driven. Statement of the principle in insufficient - it should also point the way toward a unified modelling approach that leads to prediction of the self-organised state and some relevant properties.
The first aim here is to propose a more general principle that applies beyond energy considerations. The second aim is to use the principle to formulate a framework of modelling SO in general. For didactic reasons, the principle is illustrated initially in the context of a specific example - driven two-dimensional (2D) granular systems. Then, a statistical-mechanics-based model is formulated for these systems and its extension to three-dimensional systems is outlined. A discussion how the principle applies more generally follows and it pave the way for a corresponding statistical mechanics formalism of SO in general. The application of the general formulation to a simple example of SO, where energy plays no role, is detailed in the supplemental material. In the concluding discussion, similarities to, and differences from, thermal systems in equilibrium are presented. These substantiate the generality of the principle and, more importantly, the parallels provide insight into statistical mechanics modelling of SO in other contexts. Future extensions are also proposed.
II. SO signatures in 2D granular systems
In steady states of out-of-equilibrium thermal, systems, the invested driving energy is balanced against internal dissipation. The noise in these processes is generated both by the dynamics and, depending on magnitude, also by thermal fluctuations. To illustrate the thesis here, I begin with clarifying the principle in the context of the specific example of the emergence of SO in quasi-static dynamics of driven 2D granular systems. In recent experiments and simulations, the driving methods were simple shear, pure shear, and both isotropic and anisotropic compression. The term quasi-static means that any dynamic process is slower than the response time of stress equilibration, which also means that stopping at any moment, the resulting structural configuration of the assembly of particles is in mechanical equilibrium. It is these configurations that display signatures of SO. Almost all these signatures were seen by studying one structural element - the cells. These are the smallest possible voids in the assembly, which are surrounded by particles in contact. The first signature is that distributions of several structural quantities collapse onto a master forms, once simply scaled by their means. Such a collapse was observed for many (up to 120!) different systems and different packing protocols MaBl14 . Another intriguing signature involves the evolution of cell orders, where an order is defined as the number of particles in contact around a cell. A cell orders change when contacts are made and broken and these processes resemble chemical reactions. It has been found both numerically and experimentally that the dynamics evolve toward steady states that satisfy detailed balance Waetal20 ; Suetal25 . Another observation, providing a clue to the origin of the SO, was of a competition between the stability of cells and their configurational entropy MaBl17 ; Suetal20 . The smoking gun, however, has been observations that, in a number of systems and protocols, there are strong correlations across the systems between the local principal axes of cell shapes and cell stresses MaBl17 ; MaBl21 ; Jietal25 . This strong local alignment demonstrates clearly that the local structure and local stress self-organise cooperatively, a feature that has to be included in any potential model of SO.
These observations offer insight into the principle that underlies the SO in these 2D granular dynamics. The driving and the boundary conditions, generate an internal stress field that fluctuates spatially and temporally. If a cell cannot support the local stress it disintegrates, namely, at least one of the contacts of the particles surrounding it breaks. A system-wide configuration consists of a snapshot of all its cell configurations. The SO phenomenon arises when the occurrence probabilities of a minute subset of cell shapes, orientations, and stresses configurations dominates the observations. The reason for this dominance is that these configurations of cell shapes and stresses are much more stable than the rest and therefore survive much longer. Because the lifetimes of all other configurations are too short they do not feature significantly in a typical state. In the granular systems there is a local coupling between the cell shape and stress, when unstable cells either rearrange or disintegrate, the local stress configuration around them resets. Thus, local stresses persist for longer around more stable cells and stability is correlated with the large principal stress aligning with the cell’s major axis. This interpretation is consistent with all the numerical and experimental observations and, in particular, it is almost surely the reason for the strong system-wide stress-shape orientational correlations.
III. SO statistical mechanics in granular systems
With occurrence probabilities of configurations playing a key role in SO emergence in these dynamics, a natural theoretical model for it is statistical mechanics. To formulate it, several definitions need to be introduced. A cell surrounded by $k$ particles in (force-carrying) contact is a cell of order $k$, or $k$-cell. The cell order distribution across the system is discrete, with $Q_{k}$ ($k=3,4,...,K$) the fractions of $k$-cells out of all $N_{c}$ cells. Cell stability depends on two ’attributes’: its shape and stress. To quantify the former, cells are approximated as ellipses Jietal25 . A cell ellipse has three degrees of freedom (DOFs): its major and minor axes lengths, $a_{c}$ and $b_{c}$, respectively, and the orientation of the major axis relative to a fixed frame of reference, $\theta_{c}$. Cell stresses are defined, using an appropriately weighted sum over the stresses of the particles surrounding it Bletal25 ; SM , as detailed in the supplemental material SM . The cell stress also has three degrees of freedom: the large and small principal stresses, $\sigma_{c1}$ and $\sigma_{c2}$, respectively, and the orientation of the large principal stress relative to the same fixed frame of reference, $\phi_{c}$.
At the heart of any statistical mechanical formalism is a partition function, which accounts for the occurrence probability of each system-wide configuration. In this approach, cells are regarded as quasi-particles, a configuration’s occurrence probability should increase with its stability, which would also increase its survivability. A cell’s stability depends on both its shape and its stress. The smaller the difference between the its ellipse axes lengths and its principal stresses the stabler it is. Cell stability plays a central role and stability measures need to be quantified. A standard measure of stability is VMi13
$h_{c}\equiv\frac{\sigma_{1,c}-\sigma_{2,c}}{\sigma_{1,c}+\sigma_{2,c}}\ .$ (1)
For simplicity, a similar form is assigned to the shape stability,
$f_{c}\equiv\frac{a_{c}-b_{c}}{a_{c}+b_{c}}\ .$ (2)
Next, a survivability function is required, whose value determines the configurations occurrence probabilities. It should include the stability measures, as well as any other constraints on the system. It’s main aim is to quantify the stability, and thence survivability, of system-wide configurations. The observations of function collapses and correlations between stress and shapes further suggest coupling between $\theta_{c}$ and $\phi_{c}$ at the cell level, which the survivability function should include. Additionally, it has been found Jietal25 that the most probable values of the stability measures are finite, which is another piece of information that needs to be accommodated in the survivability function. In principle, it could also include correlations between the DOFs nearest-neighbour cells, although no such correlations have been observed in the experimental and numerical systems Waetal20 ; Suetal25 . The survivability function should decreases as cell stability increases, which would make it analogous to the free energy in thermal systems. Finally, the partition function should also be grand-canonical because the number of cells, $N_{c}$, fluctuates. All these specifications are satisfied by
$Z=\prod\limits_{k=1}^{K}\left[\left(\sum_{c\in k}^{Q_{k}N_{c}}e^{-F_{c}}\right)e^{-\mu Q_{k}N_{c}}\right]\ ,$ (3)
with a cell survivability function
$F_{c}=$ $\ J_{1}f_{c}(k)+J_{2}h_{c}(k)\times$
$+J_{3}\left[\theta_{c}(k)-\phi_{c}(k)\right]^{2}+J_{c^{\prime}c}\psi_{c^{\prime}c}^{2}\ .$ (4)
In this expression, $\mu$, $J_{1}$, and $J_{2}$ are Lagrange multipliers that reflect the effect of the dynamics-generated noise on the number of cells, the cell shape stability, and the cell stress stability, respectively. $f_{0}(k)$ and $h_{0}(k)$ are the most probable values of these measures, found in Jietal25 . The last term represents the cell-cell interaction, with $\psi_{c^{\prime}c}=\left(\theta_{c}-\theta_{c^{\prime}}\right)^{2}$ the difference between the orientations of neighboring cells $c$ and $c^{\prime}$. More interaction terms between other DOFs of neighbouring cells can be added to $F_{c}$, if necessary.
As the DOFs are continuous, it makes sense to convert the summation in (3) into integrals of ‘densities of states’ over the phase space of DOFs. Denoting the cell DOFs by
$\left\{X\right\}=\left\{a,b,\theta,\sigma_{1},\sigma_{2},\phi\right\}\ ,$
the densities of states provide the occurrence probabilities of the DOFs and/or the cells stability measures. For the 2D granular systems, some of these distributions are known from modelling and observations. Specifically, $\theta_{c}$ is isotropic on the cell level and $\phi_{c}$ is strongly correlated with it Jietal25 . Thus, these two variable can be replaced by one variable, $\Delta\equiv\theta-\phi$. This variable is normally distributed almost identically around zero for all $k$, $P(\Delta)\sim e^{-\Delta^{2}/\left(2\delta^{2}\right)}$. In the absence of other information, the distributions of the stability measures $h$ and $f$ can be obtained from the distributions of the DOFs on which they depend. However, it has been shown that, in these systems, the conditional distributions of $h_{c}$ given $k$, $P_{h}(h|k)$, when scaled by the mean for each cell order, $\bar{h}(k)$, collapse onto an identical Weibull form, $W\left[\hat{h}=h/\bar{h}(k)\right]$, that is independent of $k$ Jietal25 . The observations that no cell-cell spatial correlations exist in two shear experiments Waetal20 ; Suetal25 , simplifies the analysis, $J_{c^{\prime}c}=0$, at least for these systems. This then becomes an ‘independent cells model’, . To illustrate the use of the formalism, assume for simplicity that the conditional densities of states of $f(k)$ also collapse onto a master form under a similar scaling. An explicit calculation of the partition function, detailed in the supplemental material SM , yields
[…]
$-\mu-\frac{1}{2}\ln{\left(1+2\delta^{2}J_{3}\right)}\Bigg{\}}N_{c}\ .$ (5)
This expression can be used to obtain expectation values of quantities, which depend on the DOFs and the stability measures, as functions of the noise parameters and coupling constants in the survivability function,
[…]
$P\left(\{X\}\right)d^{6}\left\{X\right\}\Bigg{\}}^{Q_{k}N_{c}}\ .$ (6)
E.g., analogously to traditional statistical mechanics, one can readily derive
$\langle f\rangle=\sum\limits_{k=3}^{K}Q_{k}\bar{f}(k)=-\frac{\partial\ln{Z}}{\partial J_{1}}$ (7)
$\langle h\rangle=\sum\limits_{k=3}^{K}Q_{k}\bar{h}(k)=-\frac{\partial\ln{Z}}{\partial J_{2}}\ .$ (8)
The extension of this formalism to three-dimensional (3D) granular SO is straightforward in principle. As in 2D, cells are the smallest polyhedral voids surrounded by particles in contact. They can be approximated as ellipsoids and their stresses defined as weighted sum of the stresses of their surrounding particles. This would give 10 DOFs per cell (three axes and two angles for each of the shape and stress ellipsoids). Similar ellipsoid and stress stability measures can be defined and $F_{c}$ can be a straightforward 3D version of (4). This, in principle, provides the partition function. However, its evaluation may be more difficult than in 3D, especially because the more complicated cells classification into families. In both 2D and 3D, a cell is the smallest void surrounded by particles in contact. This makes for polygons in 2D and polyhedra in 3D. While a polygon can be classified only by its order, $k$, classification of polyhedra requires more parameters: the number of its faces and the order of each polygonal face. Specifically, a family of cell polyhedra is classified by its $n_{p}$ vertices (contacts), $k_{3}$ triangular faces, $k_{4}$ rectangular faces, and so on up to the highest-order faces, $k_{t}$. Members of the same family have an identical series of such numbers. However, this is a quantitative, rather than qualitative, complication. It just means that the product over $k$ in (3) needs to be replaced by several products, while the integrand remains the same. This extension is under investigation currently, but it is downstream from the aim here and will be presented elsewhere.
IV. The principle and its generalisation
Emergence of SO in the above example is governed by a clear principle. Cells that are better aligned with the local principal stress direction are more stable and survive longer the fluctuating local stress. In turn, stress fluctuations result from cell rearrangements and disintegration. Thus, evolving cooperatively lengthens the lifetime of both cell stress and structure configurations. The longer local configurations survive the more frequently observed are system-wide states comprising these local configurations. This interpretation is supported by numerical and experimental observations in disordered disc assemblies. It was found that the mean area of $k$-cells, when normalised by the area of the regular $k$-polygon of the same edge length, first decreases with $k$ and then it increases again MaBl17 ; Suetal20 . The initial decrease is the result of an increasing number of elongated cell configurations, which are stable at low cell orders but become increasingly unstable as $k$ increases. This means that stress and structure are coupled and cell stability limits configurational entropy.
This principle can be extended beyond granular systems. Driving a disordered system generates noise even if thermal fluctuations are negligible. Such systems are also constrained by the environment through boundary conditions and interactions with other dynamic systems, as well as internally by the nature of interactions between their individual components. SO emerges when only a minute subset of system configurations, its microstates, survive the noise sufficiently long to dominate experimental observations. In contrast to these, the bulk of system-wide configurations disintegrate faster than observational timescales and therefore are not typical. Different driving methods may give rise to SOs of different characteristics because it is the noise that determines which configurations survive and the noise is generated by the driving.
V. Extension of the modelling framework
The extension to more general modelling of SO is as follows. (1) define quasi-particles (the cells); (2) identify the quasi-particles’ DOFs (the cell ellipse axes lengths and orientations, principal stresses and orientation); (3) identify quasi-particle stability measures that depend on the DOFs and whose values decrease with increasing quasi-particle stability (the cell stresses and shapes); (4) construct a survivability function that depends on all the quasi-particles’ stability measures and decreases monotonically as their stability increases ($F_{c}$ in eq. (4)). The value of the survivability function at each system-wide configuration is determined by the configuration’s stability, which, in turn, determines the configuration’s occurrence probability and therefore its lifetime.
Consider then a system of $N$ elements, structural or otherwise, which show signatures of SO, and define these as quasi-particles, $q=1,2,...,N$. Associated with each, are $M$ properties, indexed $m=1,2,...,M$. These properties self-organise locally for every quasi-particle across the system. Property $m$ of quasi-particle $q$ depends on $A_{m}$ of its DOFs, indexed $a_{q1}^{m},a_{q2}^{m},...,a_{qA_{m}}^{m}$. The modelling of this system requires identification of $f^{m}_{q}$ stability measure of each property $m$ of quasi-particle $q$. The quasi-particle survivability function, $F_{q}$, depends on these measure, as well as on other DOFs. It can also involve interactions between DOFs of the same quasi-particles, as well as different ones. For example, a generalisation of (4) would be
$F_{q}=$ $\sum\limits_{m=1}^{M}\sum\limits_{j=1}^{A_{m}}J^{m}_{j}\left(a^{m}_{qj}-a^{m}_{qj,0}\right)^{2}+$
$\sum\limits_{q^{\prime}}\sum\limits_{m,n=1}^{M}\sum\limits_{j=1}^{A_{m}}\sum\limits_{k=1}^{A_{n}}\lambda^{m,n}_{qj,q^{\prime}k}\left(a^{m}_{qj}-a^{n}_{q^{\prime}k}\right)^{2}\ ,$ (9)
where: $a^{m}_{qj,0}$ is the mean value of $a^{m}_{qj}$ at perfect SO; $J^{m}_{j}$ a Lagrange multiplier-like parameter that weights the significance of such a fluctuation in terms of occurrence probability; $q^{\prime}$ are all the quasi-particles that interact with $q$, e.g., its nearest neighbours; and $\lambda^{m,n}_{qj,q^{\prime}k}$ the coupling between DOF $a^{m}_{qj}$ and $a^{n}_{q^{\prime}k}$. The phase space is spanned by all the DOFs, whose number is
${\cal{N}}_{DoF}=\sum\limits_{q=1}^{N}\left\{\sum_{m=1}^{M}\left(\sum_{j=1}^{A_{m}}a_{qj}^{m}\right)\right\}\ .$
As the DOFs are usually continuous, the general canonical partition function is then
$Z=\prod\limits_{q=1}^{N}\left\{\int e^{-F_{q}-\mu}\left[\prod\limits_{m=1}^{M}\left(\prod\limits_{j=1}^{A_{m}}da^{m}_{j}\right)\right]\right\}\ .$ (10)
VI. Example application of the formalism
To illustrate the application of this formalism, consider the phenomenon of self-organised laning He71 ; HeMo95 ; Chetal00 . Laning refers to the spontaneous formation of lanes as agents (pedestrians, cars, active agents, etc.) move in a relatively crowded spaces. It is the result of agents aiming to move in a particular direction or reach a particular destination without bumping too often and too violently into others. While collisions, and their avoidance, are local, ordered system-wide laning often emerges. This SO makes the flow of agents more efficient, e.g., by minimising the time spent in the crowd before reaching the destination. In a simple version of this problem, $N$ agents move in one direction along a long strip of width that allows only few agents to move side by side. They can move at one of two speeds $v_{1}$ and $v_{2}>v_{1}$ and at any angle relative to the forward direction, $-\pi/2<\theta<\pi/2$. Moving at $\theta=0$ is fastest, but can be hampered by collisions with more slowly-moving agents. Attempting to overtake those by moving sideways may incur collisions with other agents. Here, the agents are the quasi-particles and $q=1,2,...,N$ and each agent has two DOFs,the speed and direction of movement. For each agent, the most stable mode of movement is by avoiding collisions, achievable with the stability measure,
$f_{q}=\lambda\sum\limits_{q^{\prime}\Rightarrow q}\left(\vec{v}_{q}-\vec{v}_{q^{\prime}}\right)^{2}\ .$ (11)
where $q^{\prime}\Rightarrow q$ denotes the indices of the agents nearest to $q$ (identifiable, e.g., using Voronoi tessellation). The objective of reaching the destination efficiently is a constraint that also needs to be quantified in the survivability function and this is achievable by encouraging movement in the forward direction. Thus, a natural survivability function is
$F_{q}=J\sin^{2}{\theta_{q}}+\lambda\sum\limits_{q^{\prime}\Rightarrow q}\left(\vec{v}_{q}-\vec{v}_{q^{\prime}}\right)^{2}\ .$ (12)
Further constraints and specifications can be readily introduced by adding appropriate terms to the survivability function. The corresponding partition function is then
$Z=$ $\int e^{-\sum\limits_{q=1}^{N}\left[J\sin^{2}{\theta_{q}}+\lambda\sum\limits_{q^{\prime}\Rightarrow q}\left(\vec{v}_{q}-\vec{v}_{q^{\prime}}\right)^{2}\right]}$
×$\times\prod\limits_{q=1}^{N}\Big{\{}P\left(v_{q},\theta_{q}\right)dv_{q}d\theta_{q}\Big{\}}\ ,$ (13)
with, say, $P\left(v_{q},\theta_{q}\right)=\left[p\delta(v-v_{1})+(1-p)\delta(v-v_{2})\right]/\pi$ the density of states. It is clear that, by increasing $J$ and $\lambda$, the stable system-wide configurations are in a small subset that approximate two lanes, one moving at $v_{1}$ and the other at $v_{2}$. This partition function can be readily evaluated, but this is downstream from the aim of this example, which is only to illustrate the application of the general formalism.
VII. Concluding discussion
The principle underlying emergence of SO in out-of-equilibrium driven disordered systems has been formulated. It states that SO is the result of survivability of a minute subset of system configurations, or microstates, that stand out in their exceptional stability against the noise introduced by the driving. This principle is clearly supported by observations of SO in 2D granular systems, which evolve toward stable configurations of coordinated cell structures and stresses. Based on this principle, a statistical mechanical formalism has been constructed for SO in those 2D systems and its use outlined. Similar stability measures have been identified in 3D granular systems, which makes possible a straightforward extension to a similar model and its outline has been discussed in detail. The principle goes beyond granular systems - driven disordered systems always experience noisy dynamics and, when their steady states self-organise, they exclude the overwhelming majority of theoretically possible configurations which do not survive the noise. It also goes beyond the previously proposed principle of minimisation of dissipation energy Cyetal21 ; Keetal23 , which is only a special case that applies to systems where energy is the main driver of the SO. The generality of this principle makes possible an extension of the statistical mechanics approach to describe emergence of SO in wider range of systems. This extension has been formulated here explicitly and it involves identification of quasi-particles; their DOFs; stability measures, which that depend on the DOFs; and a survivability function, whose value decreases with the stability measures. The latter is used in a Boltzmann-like exponential, which determines the occurrence probability of system configurations, and a partition function is constructed. From the partition function, all expectation values of the self-organised state can be derived. This complements the idea in Keetal23 of the system exploring the configuration space in that it parks on configurations that minimise locally the survivability function. To illustrate the general formalism, it is applied to construct a statistical mechanics model for a very simple version of the laning problem.
There are parallels between this formulation and its analogue in thermal systems, with system-wide configurations being the microstates and the stability measures paralleling the particles’ energies and potentials. The thermal survivability function is then the free energy, whose lowering increases a microstate’s occurrence probability and thence its survivability. Like the free energy, the survivability function depends on the quasi-particles’ DOFs. With this understanding on board, equilibrium states can also be viewed as self-organised. An example is solidification under temperature lowering, which is analogous to increasing $J_{i}$ in (4). the overwhelming number of microstates with very energetic particles are strongly unstable and only the minute crystalline microstates can be observed. There are, however, differences. The energy in thermal systems is a constant of the motion, from which dynamic equations for the DOFs can be derived. The survivability function only describes stability in out-of-equilibrium systems and, as far as this author knows, cannot play such a role. Additionally, thermal systems incur indigenous thermal fluctuations while in a-thermal systems it is the driving that gives rise to noise. Nevertheless, the similarities to traditional statistical mechanics are useful as guidelines for constructing general models of SO.
One can argue that the above principle also applies to the SO that is evident in the evolution of biological species. Those systems are driven by many constraints: resources, predators, environmental, etc., which they must adapt to and survive. The dynamics are then extremely noisy because the constraints magnitudes and even number can fluctuate wildly. The principle of the survival of the fittest is exactly the same, with only individual elements of the systems that can survive these fluctuations, be it cells or individuals organisms, are realised. While this may mean that, in principle, it could be possible to formulate a statistical mechanical model for species evolution, it is a daunting task due to the enormous number of potential DOFs and stability measures.
Acknowledgement:
Discussions with J.L. England and D.I. Goldman are gratefully acknowledged.